OpenStudy (anonymous):

use the binomial expansion of (p+q)^n to calculate n=6, p=0.5 binomial distribution

OpenStudy (dan815):

okay

OpenStudy (dan815):

p+q=1 p=0.5 q=0.5 so we just got 1^n = 1^6=1

OpenStudy (dan815):

no need to really expand in this case lol

OpenStudy (anonymous):

dam! thank you sooooooooooooooooooooo much!

OpenStudy (dan815):

you're welcome :)

OpenStudy (dan815):

if this p+q is not equal to 1 though, you should refer to the pascal triangle to expand fast

OpenStudy (dan815):